\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
% \MRhref is called by the amsart/book/proc definition of \MR.
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}
\begin{thebibliography}{10}

\bibitem{cupdlp}
\emph{{cuPDLP}}, \url{https://github.com/jinwen-yang/cuPDLP.jl}, accessed September 2025.

\bibitem{cupdlp-c}
\emph{{cuPDLP-C}}, \url{https://github.com/COPT-Public/cuPDLP-C}, Version 0.4.1, accessed September 2025.

\bibitem{hprlp}
\emph{{HPR-LP}}, \url{https://github.com/PolyU-IOR/HPR-LP}, 2025, Version 0.1.0, accessed August 2025.

\bibitem{applegate2021practical}
David Applegate, Mateo D{\'\i}az, Oliver Hinder, Haihao Lu, Miles Lubin, Brendan O'Donoghue, and Warren Schudy, \emph{Practical large-scale linear programming using primal-dual hybrid gradient}, Advances in Neural Information Processing Systems \textbf{34} (2021), 20243--20257.

\bibitem{applegate2025pdlp}
\bysame, \emph{Pdlp: A practical first-order method for large-scale linear programming}, arXiv preprint arXiv:2501.07018 (2025).

\bibitem{applegate2023faster}
David Applegate, Oliver Hinder, Haihao Lu, and Miles Lubin, \emph{Faster first-order primal-dual methods for linear programming using restarts and sharpness}, Mathematical Programming \textbf{201} (2023), no.~1, 133--184.

\bibitem{bertsimas1997introduction}
Dimitris Bertsimas and John~N Tsitsiklis, \emph{Introduction to linear optimization}, vol.~6, Athena Scientific Belmont, MA, 1997.

\bibitem{chambolle2011first}
Antonin Chambolle and Thomas Pock, \emph{A first-order primal-dual algorithm for convex problems with applications to imaging}, Journal of mathematical imaging and vision \textbf{40} (2011), 120--145.

\bibitem{chen2024hpr}
Kaihuang Chen, Defeng Sun, Yancheng Yuan, Guojun Zhang, and Xinyuan Zhao, \emph{Hpr-lp: An implementation of an hpr method for solving linear programming}, arXiv preprint arXiv:2408.12179 (2024).

\bibitem{dantzig1990origins}
George Dantzig, \emph{Origins of the simplex method}, A history of scientific computing, 1990, pp.~141--151.

\bibitem{dantzig1948programming}
George~B Dantzig, \emph{Programming in a linear structure}, Washington, DC (1948).

\bibitem{gleixner2021miplib}
Ambros Gleixner, Gregor Hendel, Gerald Gamrath, Tobias Achterberg, Michael Bastubbe, Timo Berthold, Philipp~M. Christophel, Kati Jarck, Thorsten Koch, Jeff Linderoth, Marco L\"ubbecke, Hans~D. Mittelmann, Derya Ozyurt, Ted~K. Ralphs, Domenico Salvagnin, and Yuji Shinano, \emph{{MIPLIB 2017: Data-Driven Compilation of the 6th Mixed-Integer Programming Library}}, Mathematical Programming Computation (2021).

\bibitem{gurobi}
{Gurobi Optimization, LLC}, \emph{{Gurobi Optimizer Reference Manual}}, 2024.

\bibitem{karmarkar1984new}
Narendra Karmarkar, \emph{A new polynomial-time algorithm for linear programming}, Proceedings of the sixteenth annual ACM symposium on Theory of computing, 1984, pp.~302--311.

\bibitem{lu2023cupdlp}
Haihao Lu and Jinwen Yang, \emph{cupdlp. jl: A gpu implementation of restarted primal-dual hybrid gradient for linear programming in julia}, arXiv preprint arXiv:2311.12180 (2023).

\bibitem{lu2024restarted}
\bysame, \emph{Restarted halpern pdhg for linear programming}, arXiv preprint arXiv:2407.16144 (2024).

\bibitem{lu2023cupdlpc}
Haihao Lu, Jinwen Yang, Haodong Hu, Qi~Huangfu, Jinsong Liu, Tianhao Liu, Yinyu Ye, Chuwen Zhang, and Dongdong Ge, \emph{cupdlp-c: A strengthened implementation of cupdlp for linear programming by c language}, arXiv preprint arXiv:2312.14832 (2023).

\bibitem{blog}
Vahab Mirrokni, \emph{Google research, 2022 \& beyond: Algorithmic advances}, \url{https://ai.googleblog.com/2023/02/google-research-2022-beyond-algorithmic.html}, 2023-02-10.

\bibitem{mittelmannbenchmark}
Hans Mittelmann, \emph{Decision tree for optimization software}, \url{http://plato.asu.edu/guide.html}, 2021.

\bibitem{nesterov2018lectures}
Yurii Nesterov et~al., \emph{Lectures on convex optimization}, vol. 137, Springer, 2018.

\bibitem{sun2025accelerating}
Defeng Sun, Yancheng Yuan, Guojun Zhang, and Xinyuan Zhao, \emph{Accelerating preconditioned admm via degenerate proximal point mappings}, SIAM Journal on Optimization \textbf{35} (2025), no.~2, 1165--1193.

\bibitem{wright1997primal}
Stephen Wright, \emph{Primal-dual interior-point methods}, SIAM, 1997.

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{definitions}
\usepackage{multirow}

\newcommand{\cupdlpx}{\texttt{cuPDLPx} }
\newcommand{\cupdlp}{\texttt{cuPDLP} }

\newcommand{\lc}{\ell_c}
\newcommand{\uc}{u_c}
\newcommand{\lv}{\ell_v}
\newcommand{\uv}{u_v}
\newcommand{\cR}{\mathcal{R}}
\newcommand{\cY}{\mathcal{Y}}

\newcommand{\cX}{\mathcal{X}}
\newcommand{\cS}{\mathcal{S}}
\newcommand{\R}{{\mathbb R}}

\title{cuPDLPx: A Further Enhanced GPU-Based First-Order Solver for Linear Programming}
\author{Haihao Lu\thanks{MIT, Sloan School of Management (haihao@mit.edu).} \and Zedong Peng\thanks{MIT, Sloan School of Management (zdpeng@mit.edu).} \and Jinwen Yang\thanks{University of Chicago, Department of Statistics (jinweny@uchicago.edu).}}
\date{}

\begin{document}
\maketitle

\begin{abstract}
    We introduce \cupdlpx, a further enhanced GPU-based first-order solver for linear programming. Building on the recently developed restarted Halpern PDHG for LP \cite{lu2024restarted}, \cupdlpx incorporates a number of new techniques, including a new restart criterion and a PID-controlled primal weight update. These improvements are carefully tailored for GPU architectures and deliver substantial computational gains. Across benchmark datasets, \cupdlpx achieves 2.5×–5× speedups on MIPLIB LP relaxations and 3×–6.8× on Mittelmann’s benchmark set, with particularly strong improvements in high-accuracy and presolve-enabled settings. The solver is publicly available at \url{https://github.com/MIT-Lu-Lab/cuPDLPx}.
    
\end{abstract}

\section{Introduction}
Linear programming (LP) is a central class of mathematical optimization problems due to its broad applicability, mathematical elegance, and computational efficiency~\cite{bertsimas1997introduction}. 
The two classic algorithmic frameworks for solving LPs are the simplex method~\cite{dantzig1948programming,dantzig1990origins} and interior-point methods (IPMs)~\cite{karmarkar1984new,wright1997primal}, both of which have shaped decades of theoretical development and practical deployment. The simplex method, introduced by Dantzig in the 1940s~\cite{dantzig1948programming}, proceeds by traversing the vertices of the feasible polyhedron and has demonstrated remarkable efficiency in practice despite its exponential worst-case complexity. Its numerical robustness and interpretability have made it a mainstay in commercial solvers. In contrast, IPMs operate within the interior of the feasible region, iteratively following a central path toward optimality. Since the seminal work of Karmarkar in the 1980s~\cite{karmarkar1984new}, IPMs have offered polynomial-time theoretical guarantees and have been extended to handle large and structured problems with high numerical precision. Modern LP solvers often integrate both approaches, exploiting the simplex method’s flexibility and warm-start capability alongside the theoretical convergence guarantees and scalability of IPMs. These methods remain the foundation of the state-of-the-art CPU-based LP solvers, underpinning their ability to deliver high-accuracy solutions across diverse problem instances.

Recently, first-order methods (FOMs)~\cite{nesterov2018lectures} have emerged as a compelling alternative for solving large-scale LPs, offering a new paradigm that complements the traditional dominance of simplex and interior-point methods. Unlike simplex and interior-point methods, which rely on matrix factorizations and sequential computations, FOMs operate using simple iterative updates, primarily matrix-vector multiplications, making them particularly attractive for large-scale LPs. 
A notable breakthrough in this space is PDLP~\cite{applegate2021practical,applegate2025pdlp}, a solver based on the restarted average primal-dual hybrid gradient method (PDHG)~\cite{applegate2023faster, chambolle2011first} and tailored specifically for LP. By incorporating various practical enhancements, PDLP achieves significantly improved numerical stability and convergence in practice, positioning it as a practical and scalable solver for large-scale LPs within the FOM framework.

Building on the CPU-based PDLP, recent efforts have turned toward leveraging graphics processing units (GPUs) to further accelerate LP solvers. GPUs offer massive parallelism and high memory bandwidth, making them particularly well-suited for the core computational kernels of first-order methods, namely, sparse matrix-vector multiplications and vector operations. As LP problem sizes continue to grow, sometimes even reaching billions of variables and constraints in modern applications~\cite{blog}, harnessing GPU architectures becomes increasingly critical for achieving scalable performance of first-order methods. This shift has led to the development of a new generation of GPU-accelerated LP solvers that combine algorithmic efficiency with hardware-aware implementation strategies. 

Particularly, \cupdlp\cite{lu2023cupdlp,lu2023cupdlpc} is a GPU-accelerated solver that extends PDLP by adapting its first-order framework to modern GPU architectures, with several algorithmic changes to make it more suitable for GPU architectures. The initial implementation, developed in Julia and referred to as \texttt{cuPDLP}~\cite{lu2023cupdlp}, offloads key linear algebra operations, such as sparse matrix-vector multiplications, to the GPU and incorporates GPU-friendly heuristics, achieving notable speedups on medium-to-large-scale LP instances. Serving as both a practical solver and a research platform, \texttt{cuPDLP} illustrates how careful alignment between algorithm design and hardware capabilities can yield substantial performance improvements. A subsequent C implementation, \texttt{cuPDLP-C}~\cite{lu2023cupdlpc}, was developed to facilitate integration with production-scale computing environments. These advances in GPU-based LP solvers have attracted strong interest from both optimization software companies and technology firms, and have influenced the design of several commercial solvers, including Gurobi, COPT, FICO Xpress and NVIDIA’s cuOpt.

More recently, a distinct approach based on the Halpern Peaceman–Rachford (HPR) method~\cite{sun2025accelerating} was proposed. Building on this framework, HPR-LP~\cite{chen2024hpr} was developed as a GPU-based LP solver, and numerical results on standard benchmarks demonstrate superior performance compared to \texttt{cuPDLP}, especially in obtaining high-accuracy solutions.

These recent GPU-based solvers demonstrate the growing interest and practical potential of first-order methods when carefully adapted to modern hardware. Building on these foundations, we present \cupdlpx, a further enhanced GPU-accelerated linear programming solver. Compared to its predecessor \texttt{cuPDLP}, \cupdlpx introduces several fundamental algorithmic changes that are primarily motivated by recent theoretical developments. 
\begin{itemize}
    \item While \texttt{cuPDLP} is based on the restarted averaged PDHG (raPDHG) method~\cite{applegate2023faster}, \cupdlpx adopts the restarted Halpern PDHG (rHPDHG), a refinement inspired by recent theoretical insights\cite{lu2024restarted}. This change allows the algorithm to take more aggressive steps, resulting in stronger empirical performance. The observed numerical improvements of \cupdlpx over \texttt{cuPDLP} highlight the practical advantages of the rHPDHG framework.

    \item In addition, \cupdlpx explores a distinct set of enhancements and heuristics that further accelerate performance. These include a constant step-size rule, a restart condition aligned with the theoretical guarantees of rHPDHG~\cite{lu2024restarted}, and a novel PID controller for updating the primal weight.

\end{itemize}
In addition to these algorithmic advances, \cupdlpx is developed in C, providing a lightweight and highly efficient implementation that complements the earlier Julia prototype. These improvements together translate into strong empirical performance of \cupdlpx. On the MIPLIB relaxations, \cupdlpx achieves a 2.5×–5× speedup over \texttt{cuPDLP}, while on Mittelmann’s benchmark set, the gains are even larger, reaching 3×–6.8× speedups across different tolerances and presolve settings.

The solver is publicly available at \url{https://github.com/MIT-Lu-Lab/cuPDLPx}.

\subsection{Paper organization}
Section \ref{sec:prelim} briefly introduces the form of LP to solve and the base algorithm, vanilla PDHG. In Section \ref{sec:enhancement}, several algorithmic enhancements on top of \cupdlp are proposed to accelerate convergence. The numerical comparisons between \cupdlpx and \cupdlp are presented in Section \ref{sec:numerical}.

\subsection{Notations}
For a symmetric positive semidefinite matrix $M \in \mathbb{R}^{n \times n}$, the {\( M \)-norm} of a vector \( x \in \mathbb{R}^n \) is defined as $\|x\|_M := \sqrt{x^\top M x}$. The orthogonal projector \( \operatorname{proj}_C(x) \) denotes the Euclidean projection of a point \( x \in \mathbb{R}^n \) onto a closed convex set \( C \subseteq \mathbb{R}^n\), i.e., $\operatorname{proj}_C(x) := \arg\min_{y \in C} \|x - y\|_2$. For a set \( C \subseteq \mathbb{R}^n \), we define \( -C := \{-x : x \in C\} \) as the {reflection of \( C \) about the origin}. The \((i,j)\)-th entry for a matrix \( M \in \mathbb{R}^{m \times n} \) is denoted by \( (M)_{ij} \). Given two vectors \( \ell \in (\mathbb{R} \cup \{-\infty\})^n \) and \( u \in (\mathbb{R} \cup \{\infty\})^n \) satisfying \( \ell \leq u \), define the mapping \( p: \mathbb{R}^n \times (\mathbb{R} \cup \{-\infty\})^n \times (\mathbb{R} \cup \{\infty\})^n \to \mathbb{R} \cup \{\infty\} \) given by $p(y; \ell, u) := u^\top y^+ - \ell^\top y^-$.

\section{Preliminaries}\label{sec:prelim}
In this section, we introduce the LP form that \cupdlpx solves, followed by discussions on vanilla PDHG for solving LPs.

\subsection{Linear programming}
\cupdlpx solves LP with the following primal-dual form:
\begin{equation}\label{eq:lp}
    \begin{aligned}[c]
    \min_{x\in \cX}~~ &~ c^\top x & \\
\text{subject to:}~~ &~ Ax\in \cS \ ,
    \end{aligned}
    \qquad\qquad\qquad
    \begin{aligned}[c]
\max_{y\in \cY, r \in\cR} \quad &  - p(-y; \lc, \uc) - p(-r ; \lv, \uv) \\
\text{subject to:}\quad & c - A^\top y = r \ ,
    \end{aligned}
\end{equation} 
where $\cX := \{x \in \R^n : \lv \leq x \leq \uv \}$ with $\lv \in (\R \cup \{ -\infty \})^{n}$ and $\uv \in (\R \cup \{ \infty \})^{n}$, $\cS:=\{s\in\R^m: \lc\leq s \leq\uc\}$ with $\lc \in (\R \cup \{- \infty\})^m$ and $\uc \in (\R \cup \{\infty\})^m$, $A \in \R^{m \times n}$, $c \in \R^{n}$, and the sets $\cY\subseteq \R^{m}$ and $\cR \subseteq \R^{n}$ are Cartesian products with their $i$th components given by 
\begin{equation*}
\begin{aligned}
    &\cY_i := \begin{cases}
\{ 0 \} & (\lc)_i = -\infty, ~ (\uc)_i = \infty,  \\
\R^{-} & (\lc)_i = -\infty, ~ (\uc)_i \in \R ,\\
\R^{+} & (\lc)_i \in \R, ~ (\uc)_i = \infty, \\
\R & \text{otherwise}; 
\end{cases}
\quad \text{ and } \qquad
\end{aligned}
\begin{aligned}
    &\cR_i := \begin{cases}
\{ 0 \} & (\lv)_i = -\infty, ~ (\uv)_i = \infty,  \\
\R^{-} & (\lv)_i = -\infty, ~ (\uv)_i \in \R, \\
\R^{+} & (\lv)_i \in \R, ~ (\uv)_i = \infty, \\
\R & \text{otherwise};
\end{cases} \ .
\end{aligned}
\end{equation*}
Equivalently the primal-dual form of \eqref{eq:lp} is
\begin{flalign}\label{eq:lp-pd}
\max_{y \in \cY}\min_{x \in \cX}\   L(y,x) := - p(y; -\uc, -\lc) +  y^\top A x +c^\top x \ .
\end{flalign} 
This LP form is used in CPU-based PDLP~\cite{applegate2025pdlp}.
\subsection{PDHG} 
PDHG serves as the base routine of \cupdlpx to solve the primal-dual problem \eqref{eq:lp-pd}. Specifically, the update rule of PDHG on \eqref{eq:lp-pd} is given as:
\begin{equation}\label{eq:pdhg}
    \begin{aligned}
        & x^{k+1} = \operatorname{proj}_{\cX}\pran{x^k-{\tau}(c-A^\top y^k)} \\
        & y^{k+1}=y^k-\sigma A(2x^{k+1}-x^k)-\sigma\operatorname{proj}_{-\cS}\pran{\sigma ^{-1}y^k-A(2x^{k+1}-x^k)}\ , 
    \end{aligned}
\end{equation}
where $\sigma$ and $\tau$ are dual and primal stepsizes respectively. \cupdlpx further reparameterizes as $\tau=\frac{\eta}{\omega}$ and $\sigma=\eta\omega$, where $\eta$ is called the stepsize and $\omega$ is the primal weight. And the canonical norm of PDHG is defined by $\|\cdot\|_P$, where $P:=P_{\eta,\omega}=\begin{bmatrix}
    \frac{\omega}{\eta}I & A^\top\\ A & \frac{1}{\eta\omega}I
\end{bmatrix}$. For notational convenience, we define the primal-dual iterate at iteration $k$ as $z^k = \begin{bmatrix} x^k \\ y^k \end{bmatrix}$, and write $z^{k+1} = \operatorname{PDHG}(z^k)$ to denote a single PDHG update step applied to $z^k$ as defined in~\eqref{eq:pdhg}.

\section{Algorithmic enhancements on top of \cupdlp}\label{sec:enhancement}

In this section, we discuss a series of algorithmic further enhancements used in \cupdlpx to improve the efficiency and convergence behavior of the algorithm, on top of what is used in \cupdlp. These enhancements are primarily motivated by recent theoretical advances in restarted Halpern PDHG, as described in~\cite{lu2024restarted}.

At the high-level, \cupdlpx is based on restarted Halpern PDHG algorithm (in contrast to the restarted average PDHG algorithm used in \cupdlp), and includes additional techniques such as the reflected updates, adaptive restarting, stepsize selection and dynamic primal-dual weight. Each component is carefully designed and implemented to accelerate convergence and enhance stability for solving linear programming problems. We describe these contributions below.

\textbf{Halpern scheme.} The Halpern scheme, along with its reflected variant, is an enhancement to the PDHG algorithm that improves convergence properties and accelerates solver performance. Originally developed to accelerate general operator splitting methods, the Halpern scheme has been successfully adapted to linear programming with strong theoretical guarantees and improved preliminary numerical performance~\cite{lu2024restarted}. 
Halpern PDHG interpolates between the current PDHG iterate and the initial point, using a weighted average. Specifically, the update rule at iteration $k$ is:
\begin{equation*}
    z^{k+1}=\frac{k+1}{k+2}\operatorname{PDHG}(z^k)+\frac{1}{k+2}z^0\ .
\end{equation*}

As discussed in~\cite{lu2024restarted}, the Halpern and average schemes share structural similarities but differ in theoretical guarantees. Notably, the Halpern scheme supports reflection, which enables the use of larger stepsizes and often leads to faster empirical convergence, as described below.

\textbf{Reflection.} The reflection technique builds upon the Halpern scheme by applying a reflected version of the PDHG operator, $(1+\gamma)\operatorname{PDHG}-\gamma \operatorname{id}$ with reflection $\gamma \in [0,1]$, instead of the vanilla PDHG update. This is also called over-relaxation in the literature of variational inequality. The resulting update rule of the Reflected Halpern PDHG update is
\begin{equation*}
    z^{k+1}=\frac{k+1}{k+2}\pran{(1+\gamma)\operatorname{PDHG}(z^k)-\gamma z^k}+\frac{1}{k+2}z^0\ .
\end{equation*}
This reflection mechanism effectively allows for longer steps than the vanilla Halpern update and has been shown, both theoretically and empirically, to improve convergence~\cite{lu2024restarted}. It is one of the key drivers of \cupdlpx’s superior performance compared to \cupdlp.

\textbf{Adaptive restart.} Restarting is a key enhancement on first-order methods for attaining high-accuracy solutions for linear programming~\cite{applegate2023faster}. A restart is triggered when certain progress criteria are met, and the algorithm restarts from a new initial solution. In the Halpern scheme, the anchor is periodically reset to the current solution. This helps the algorithm stay focused on the neighborhood of the optimal solution, especially as the initial anchor becomes increasingly outdated during iterations. Such restart strategies have been shown to accelerate convergence rates for Halpern PDHG~\cite{lu2024restarted} in theory.

\cupdlpx adopts an adaptive restart strategy that evaluates potential restart conditions at each iteration. The core idea is to monitor a fixed-point error metric $r(z)=\|z-\operatorname{PDHG}(z)\|_P$ at solution $z$, and trigger a restart when specific decay patterns are observed. This fixed-point error metric is motivated by the recent theoretical insight~\cite{lu2024restarted}, which shows that this fixed-point error is a natural metric for the restarted Halpern PDHG. It is different from the normalized duality gap used in CPU-based PDLP~\cite{applegate2021practical,applegate2025pdlp} and the KKT error used in cuPDLP~\cite{lu2023cupdlp}.   

Three restart conditions are used:
\begin{itemize}
    \item (sufficient decay) $r(z^{n,k})\leq \beta_{\mathrm{sufficient}}r(z^{n,0})$
    \item (necessary decay + no local progress) $r(z^{n,k})\leq \beta_{\mathrm{necessary}}r(z^{n,0})$ and $r(z^{n,k})>r(z^{n,k-1})$
    \item (artificial restart) $k\geq \beta_{\mathrm{artificial}} T$, where $T$ is the total iteration.
\end{itemize}

This combined restart strategy balances aggressiveness and stability, contributing to cuPDLPx’s robust and accelerated convergence across a wide range of linear programming instances.

\textbf{Stepsize.} In contrast to an adaptive stepsize heuristic used in \texttt{cuPDLP}, \cupdlpx switches to constant stepsize $\eta=\frac{0.998}{\|A\|_2}$, where operator norm of constraint matrix $\|A\|_2$ is approximated by power iteration. The rationale behind this change is twofold. First, the reflected Halpern PDHG inherently selects a larger effective stepsize, and the adaptive stepsize strategy from~\cite{lu2023cupdlp} can become overly aggressive in certain instances, leading to occasional instability. Second, a constant stepsize eliminates the need for the sequential stepsize search, which is not perfectly-suited for parallel execution on GPUs. Thus, this modification improves both stability and parallel efficiency in cuPDLPx.

\textbf{Primal weight update.} 
The primal weight $\omega$ is a crucial parameter in balancing progress between the primal and dual spaces. Our empirical experiences have shown that convergence behavior in PDLP is highly sensitive to the choice of this weight. The intuition for adjusting the primal weight $\omega^n$ is to set it such that the weighted distances to optimality in the primal and dual domains are in the same scale, i.e., 
$\|(x^{n,t} - x^*), 0\|_{w^n} \approx \|(0, y^{n,t} - y^*)\|_{w^n}$. 
However, the optimal solution $(x^*, y^*)$ is not known during the iterations. 

To address this, \texttt{cuPDLP} estimates the primal and dual distances to optimality based on the observed movements in the previous epoch. The distance estimates are smoothed using exponential averaging and subsequently used to update the primal weight at each restart.

In \cupdlpx, we further enhance this strategy by modeling the update as a control problem and introducing a PID  controller to dynamically regulate the primal weight. 
We define the error as the gap between the primal and dual distances on a logarithmic scale:
\[
e^n = \log \left( \frac{\sqrt{w^n} \left\| x^{n,t} - x^{n,0} \right\|_2}{\frac{1}{\sqrt{w^n}} \left\| y^{n,t} - y^{n,0} \right\|_2}  \right)
\]

The primal weight is then updated according to:
\[
\log w^{n+1} = \log w^n - [K_P \cdot e^n + K_I \cdot \sum_{i=1}^{n} e^i + K_D \cdot (e^n - e^{n-1})]
\]
where $K_P$, $K_I$ and $K_D$ are the coefficients for the proportional, integral and derivative terms, respectively. The update is applied at each restart occurrence and the initial primal weight is set to $1.0$.

This PID-controlled update ensures that the primal-dual balance is continuously adjusted in a stable manner, improving convergence speed and robustness across a wide range of problem instances.

\section{Numerical experiments}\label{sec:numerical}
In this section, we compare the numerical performance of \cupdlpx with \texttt{cuPDLP}. We first describe the setup of the experiments, followed by presenting the numerical results on  LP relaxations of instances from the MIPLIB 2017 collection and Mittelmann’s LP benchmark. 

\textbf{Benchmark datasets.}
We conduct numerical experiments on two widely used LP benchmark datasets: MIP Relaxations, which consists of 383 instances derived from root-node LP relaxations of MIPLIB 2017~\cite{gleixner2021miplib}, and Mittelmann’s LP benchmark dataset~\cite{mittelmannbenchmark}, which contains 49 LP instances. In particular, the 383 instances in the MIP relaxations dataset are selected based on the same filtering criteria outlined in~\cite{lu2023cupdlp}. Based on the number of non-zeros of the constraint matrices, we further categorize these 383 instances into three groups, as summarized in Table~\ref{tab:miplib-size}.
\begin{table}[ht!]
\centering
\begin{tabular}{cccc}
\hline
                    & \textbf{Small}                   & \textbf{Medium}                     & \textbf{Large}                   \\ \hline
\textbf{Number of nonzeros}  & 100K -  1M & 1M - 10M & \textgreater 10M \\
\textbf{Number of instances} & 269                     & 94                         & 20                      \\ \hline
\end{tabular}
\caption{Scale and number of instances in \texttt{MIP Relaxations}.}
\label{tab:miplib-size}
\end{table}

\textbf{Software.} 
We implement \cupdlpx in both C and Julia. The Julia version utilizes \texttt{CUDA.jl} as the interface for working with NVIDIA CUDA GPUs, while the C version directly leverages low-level CUDA libraries for improved performance. We compare the performance of  \cupdlpx against \texttt{cuPDLP} \cite{cupdlp}, \texttt{HPR-LP v0.1.0} \cite{hprlp}, \texttt{Gurobi 12.0.4} \cite{gurobi}, and \texttt{cuPDLP-C v0.4.1} \cite{cupdlp-c}.
All Julia implementations and codes are timed after pre-compilation. 
For the Gurobi experiments, we set \texttt{Method=6} to activate PDLP and \texttt{GURO\_PAR\_PDHGGPU=1} to enable GPU acceleration.

\textbf{Computing environment.} 
We use an NVIDIA H100-SXM-80GB GPU, with CUDA 12.4. The experiments of \texttt{cuPDLP}, \texttt{HPR-LP} and \cupdlpx (Julia version) are performed in Julia 1.11.5.

\textbf{Initialization.} 
\cupdlpx uses all-zero vectors as the initial starting points.

\textbf{Optimality termination criteria.} \cupdlpx terminate when the relative KKT error is no greater than the termination tolerance $\epsilon \in (0, \infty)$:
\begin{equation*}
    \begin{aligned}
    &|c^\top x + p(-y; \lc, \uc) + p(-r ; \lv, \uv)|\leq \epsilon(1+| p(-y; \lc, \uc) + p(-r ; \lv, \uv)|+|c^\top x|)\\
    &\left\|Ax-\operatorname{proj}_{[L,U]}(Ax) \right\|_2 \leq  \epsilon \pran{1+\left\|(L, U) \right\|_2 } \\
    &\left\|c-A^\top y-r\right\|_2 \leq \epsilon (1+\|c\|_2)\\
    &\left\|r-\operatorname{proj}_{\mathcal R}(r)\right\|_2  \leq \epsilon(1+\|c\|_2) \ .
    \end{aligned}
\end{equation*}
The termination criteria are checked for the original LP instance, not the preconditioned ones, so that the preconditioning does not impact the termination. In the experiment, we set $\epsilon=10^{-4}$ for moderate accuracy and $\epsilon=10^{-8}$ for high accuracy. 
For Gurobi, both absolute and relative feasibility tolerances are checked, and feasibility is considered satisfied once either tolerance is met. To align termination criteria across solvers, we fix the absolute feasibility tolerance to \texttt{GURO\_PAR\_PDHGABSTOL=1e-9}, while \texttt{GURO\_PAR\_PDHGCONVTOL} and \texttt{GURO\_PAR\_PDHGRELTOL} serve as the effective convergence criteria in Gurobi.

\textbf{Time limit.} 
We apply a time limit of 3600 seconds to small and medium-sized problems and 18000 seconds to large-scale problems for MIPLIB relaxation.
A 15000-second time limit is used for Mittelmann’s benchmark.

\textbf{Shifted geometric mean.} 
We use the shifted geometric mean of solving time to evaluate solver performance across a collection of instances. Formally, the shifted geometric mean is defined as $\left(\prod_{i=1}^n (t_i+\Delta)\right)^{1/n}-\Delta$ where $t_i$ is the solve time for the $i$-th instance. We shift by $\Delta = 10$ and denote this metric as SGM10.  If an instance is unsolved, its solving time is set to the corresponding time limit.

Tables~\ref{tab:miplib-no-presolve} and~\ref{tab:miplib-presolve} compare the performance of \texttt{cuPDLP} and \cupdlpx on 383 MIPLIB instances, with and without Gurobi presolve, respectively. Several key observations can be made:

\begin{itemize}
\item Across all four settings, \cupdlpx consistently outperforms \texttt{cuPDLP} in terms of overall numerical performance, solving more instances and achieving notable reductions in running time.
\item The performance gain of \cupdlpx is especially significant for Small and Medium instances, while the improvement is less significant for Large instances, likely due to the high variance among the relatively small number of Large instances.
\end{itemize}

Table~\ref{tab:speedup} reports the speedup of \cupdlpx over \texttt{cuPDLP}. In particular, the first row focuses on hard instances, defined as those requiring at least 10 seconds to solve using either \texttt{cuPDLP} or \cupdlpx. In various scenarios, \cupdlpx achieves a 2× to 5× speedup. The improvement is more substantial when Gurobi presolve is enabled compared to the case without presolve. The speedup is also more significant on harder instances. For example, in the high-accuracy setting with presolve, \cupdlpx achieves a 3.57× speedup on all instances and a 4.99× speedup on hard instances. These results demonstrate that the algorithmic enhancements of \cupdlpx are the primary driver of the observed 2×–5× improvements, while the C implementation yields an additional performance gain of about 20\% compared to the Julia prototype.

\begin{table}[h!]
\centering
{\small
\begin{tabular}{cccrcrcrcr}
\toprule
\multirow{2}{*}{$\epsilon$}     &    \multirow{2}{*}{\textbf{Solver}}                          & \multicolumn{2}{c}{\begin{tabular}[c]{@{}c@{}}\textbf{Small (269)} \\ (1-hour limit)\end{tabular}} & \multicolumn{2}{c}{\begin{tabular}[c]{@{}c@{}}\textbf{Medium (94)}\\ (1-hour limit)\end{tabular}}    & \multicolumn{2}{c}{\begin{tabular}[c]{@{}c@{}}\textbf{\textbf{Large (20)}}\\ (5-hour limit)\end{tabular}} & \multicolumn{2}{c}{\textbf{Total (383)}}          \\
                                                    & &\textbf{Count} & \textbf{Time}  & \textbf{Count} & \textbf{Time}  & \textbf{Count} & \textbf{Time} & \textbf{Count} & \textbf{Time}  \\ \midrule
\multirow{6}{*}{$\mathbf{10^{-4}}$}  &\multicolumn{1}{l}{\textbf{cuPDLP} (Julia)}     & 266                   & 8.81                & 91                    & 11.55                & {19}                    & 77.43 &376 &11.07           \\
&\multicolumn{1}{l}{\textbf{HPR-LP} (Julia)}          & 268 & 3.94  & 92 & 6.46   & 15 & 150.85 & 375 & 6.50 \\
&\multicolumn{1}{l}{\textbf{cuPDLPx} (Julia)}     & {269}                   & {3.50}               & {93}                    & {5.51}              & {19}                    & {62.04} & {381} & {5.24}\\
&\multicolumn{1}{l}{\textbf{cuPDLP-C}}           & 268 & 3.78  & 92 & 7.54   & 18 & 60.12  & 378 & 5.92 \\

&\multicolumn{1}{l}{\textbf{Gurobi-PDLP}}        & 267 & 4.74  & 92 & 7.67   & 17 & 103.47 & 376 & 7.14 \\
&\multicolumn{1}{l}{\textbf{cuPDLPx} (C)}     &  269                 &  2.77              &      94               &      4.62       &    19                & 57.18 & 382 & 4.39\\
\midrule
\multirow{6}{*}{$\mathbf{10^{-8}}$}  &\multicolumn{1}{l}{\textbf{cuPDLP} (Julia)}     & 260                   & 24.80               & 87                    & 36.90                & {17}                    & {208.66}  &364 &31.22             \\
&\multicolumn{1}{l}{\textbf{HPR-LP} (Julia)}          & 261 & 10.77 & 91 & 16.78  & 14 & 485.71 & 366 & 16.09 \\
&\multicolumn{1}{c}{\textbf{cuPDLPx} (Julia)}     & {264}                   & {8.94}               & {91}                    & {16.43}              & {17}                   & {194.91} & {372} & {13.27}  \\
&\multicolumn{1}{l}{\textbf{cuPDLP-C}}           & 265 & 11.72 & 87 & 20.81  & 16 & 186.44 & 368 & 16.55 \\
&\multicolumn{1}{l}{\textbf{Gurobi-PDLP}}        & 261 & 12.22 & 89 & 19.77  & 17 & 388.87 & 367 & 17.76 \\
&\multicolumn{1}{l}{\textbf{cuPDLPx} (C)}     &  266                 &   6.84             &     92                & 13.91            &  16                  &194.40  &374  &10.91 \\
\bottomrule
\end{tabular}
}
\caption{Solve time in seconds and SGM10 on instances of {MIP Relaxations} without presolve.}
\label{tab:miplib-no-presolve}
\end{table}

\begin{table}[h!]
\centering
{\small
\begin{tabular}{cccrcrcrcr}
\toprule
\multirow{2}{*}{$\epsilon$}     &   \multirow{2}{*}{\textbf{Solver}}                          & \multicolumn{2}{c}{\begin{tabular}[c]{@{}c@{}}\textbf{Small (269)} \\ (1-hour limit)\end{tabular}} & \multicolumn{2}{c}{\begin{tabular}[c]{@{}c@{}}\textbf{Medium (94)}\\ (1-hour limit)\end{tabular}}    & \multicolumn{2}{c}{\begin{tabular}[c]{@{}c@{}}\textbf{\textbf{Large (20)}}\\ (5-hour limit)\end{tabular}} & \multicolumn{2}{c}{\textbf{Total (383)}}          \\
                                                    & &\textbf{Count} & \textbf{Time}  & \textbf{Count} & \textbf{Time}  & \textbf{Count} & \textbf{Time} & \textbf{Count} & \textbf{Time}  \\ \midrule
\multirow{6}{*}{$\mathbf{10^{-4}}$}  &\multicolumn{1}{l}{\textbf{cuPDLP} (Julia)}     & 268                   & 5.11                & 92                    & 9.03                & {19}                    & 26.74 &379 &6.75          \\
&\multicolumn{1}{l}{\textbf{HPR-LP} (Julia)}          & 268 & 2.21  & 94 & 3.88   & 19 & 15.32  & 381 & 3.08 \\
&\multicolumn{1}{l}{\textbf{cuPDLPx} (Julia)} &	269 & 1.85 & 94 & 3.36 & 19 & 15.06 & 382 & 2.69 \\
&\multicolumn{1}{l}{\textbf{cuPDLP-C}}           & 268 & 1.99  & 94 & 4.25   & 19 & 13.25  & 381 & 2.95 \\
&\multicolumn{1}{l}{\textbf{Gurobi-PDLP}}        & 269 & 2.57  & 93 & 5.09   & 19 & 16.52  & 381 & 3.67 \\
&\multicolumn{1}{l}{\textbf{cuPDLPx} (C)}& 269&	1.41&	94&	2.75&	19&	14.07&	382&	2.19\\
\midrule
\multirow{6}{*}{$\mathbf{10^{-8}}$}  &\multicolumn{1}{l}{\textbf{cuPDLP} (Julia)}     & 264                   & 18.50             & 90                    & 29.40                & {19}                    & {63.68}  &373 &22.42             \\
&\multicolumn{1}{l}{\textbf{HPR-LP} (Julia)}          & 268 & 5.85  & 92 & 11.90  & 19 & 45.63  & 379 & 8.32 \\
&\multicolumn{1}{l}{\textbf{cuPDLPx} (Julia)}&	269&	5.16&	93&	11.02&	19&	46.84&	381&	7.60\\
&\multicolumn{1}{l}{\textbf{cuPDLP-C}}           & 264 & 9.28  & 89 & 16.05  & 18 & 46.70  & 371 & 11.96 \\
&\multicolumn{1}{l}{\textbf{Gurobi-PDLP}}        & 265 & 7.35  & 90 & 14.52  & 19 & 39.65  & 374 & 9.95 \\
&\multicolumn{1}{l}{\textbf{cuPDLPx} (C)}& 269&	3.98&	92&	9.60&	19&	43.09&	380&	6.28\\
\bottomrule
\end{tabular}
}
\caption{Solve time in seconds and SGM10 on instances of {MIP Relaxations} with Gurobi presolve.}
\label{tab:miplib-presolve}
\end{table}

\begin{table}[ht!]
\centering
\begin{tabular}{cccccc}
\toprule
\multirow{2}{*}{\textbf{Instance}} & \multirow{2}{*}{\textbf{Version}} 
& \multicolumn{2}{c}{\textbf{Without Presolve}} & \multicolumn{2}{c}{\textbf{With Presolve}} \\
\cmidrule(r){3-4} \cmidrule(l){5-6}
& & $10^{-4}$ & $10^{-8}$ & $10^{-4}$ & $10^{-8}$ \\
\midrule
\multirow{2}{*}{\textbf{Hard}} 
  & Julia & 3.82 & 3.09 & 4.06 & 4.14 \\
  & C     & 4.60 & 3.81 & 4.85 & 4.99 \\
\midrule
\multirow{2}{*}{\textbf{Overall}} 
  & Julia & 2.11 & 2.35 & 2.51 & 2.95 \\
  & C     & 2.52 & 2.86 & 3.08 & 3.57 \\
\bottomrule
\end{tabular}
\caption{Summarization of \texttt{cuPDLPx} speedup (Julia and C versions) over \texttt{cuPDLP}. 
The ``Hard'' rows report speedup on instances where either solver takes more than 10 seconds, 
while the ``Overall'' rows summarize speedup across all instances.}
\label{tab:speedup}
\end{table}

Tables~\ref{tab:mittelmann-no-presolve} and~\ref{tab:mittelmann-presolve} summarize results on Mittelmann’s LP benchmark set with 49 instances. Consistent with the MIPLIB experiments, \cupdlpx achieves clear speedups over \texttt{cuPDLP} across all tolerances and presolve settings. At tolerance $10^{-4}$, \cupdlpx yields about a 3×–4× speedup without presolve, which increases to 6.8× with presolve. At the tighter tolerance $10^{-8}$, the improvements remain substantial, with about a 3× speedup without presolve and over 5× with presolve.

\begin{table}[ht!]
\centering
\begin{tabular}{ccrcr}
\toprule
\multirow{2}{*}{\textbf{Solver}} 
                                & \multicolumn{2}{c}{$\mathbf{\epsilon=10^{-4}}$}      & \multicolumn{2}{c}{$\mathbf{\epsilon=10^{-8}}$}       \\ 
                                                    & \textbf{Count} & \textbf{Time} & \textbf{Count} & \textbf{Time} \\ 
                                                    \midrule
\multicolumn{1}{l}{\textbf{cuPDLP} (Julia)}     & 43                     &   67.38             &  40                   &    219.51           \\
\multicolumn{1}{l}{\textbf{HPR-LP} (Julia)} & 48 & 24.28 & 42 & 96.08 \\
\multicolumn{1}{l}{\textbf{cuPDLPx} (Julia)}     & {48}                     &   {22.78}             &  {44}                   &    {78.62}          \\
\multicolumn{1}{l}{\textbf{cuPDLP-C}} & 47 & 35.36 & 40 & 141.02 \\
\multicolumn{1}{l}{\textbf{Gurobi-PDLP}} & 44 & 32.47 & 39 & 153.94 \\
\multicolumn{1}{l}{\textbf{cuPDLPx} (C)}     &   48                &   19.11             &    44                 &   70.64          \\
\bottomrule
\end{tabular}
\caption{Solve time in seconds and SGM10 on instances of {Mittelmann's LP} benchmark set without presolve.}
\label{tab:mittelmann-no-presolve}
\end{table}

\begin{table}[ht!]
\centering
\begin{tabular}{ccrcr}
\toprule
\multirow{2}{*}{\textbf{Solver}} & \multicolumn{2}{c}{$\mathbf{\epsilon=10^{-4}}$} & \multicolumn{2}{c}{$\mathbf{\epsilon=10^{-8}}$} \\
 & \textbf{Count} & \textbf{Time} & \textbf{Count} & \textbf{Time} \\ \midrule
\multicolumn{1}{l}{\textbf{cuPDLP} (Julia)} & 46 & 48.76 & 42 & 187.22 \\
\multicolumn{1}{l}{\textbf{HPR-LP} (Julia)} & 49 & 11.85 & 47 & 53.17 \\
\multicolumn{1}{l}{\textbf{cuPDLPx} (Julia)} & 49 & 9.30  & 48 & 38.96 \\
\multicolumn{1}{l}{\textbf{cuPDLP-C}} &  47 & 19.32 & 45 & 103.82 \\
\multicolumn{1}{l}{\textbf{Gurobi-PDLP}} & 46 & 21.95 & 43 & 93.29 \\
\multicolumn{1}{l}{\textbf{cuPDLPx} (C)}     & 49 & 7.09  & 47 & 33.70 \\ \bottomrule
\end{tabular}
\caption{Solve time in seconds and SGM10 on instances of {Mittelmann's LP} benchmark set with Gurobi presolve.}
\label{tab:mittelmann-presolve}
\end{table}

\section{Conclusions}
In this paper, we present \cupdlpx, an enhanced GPU implementation of restarted Halpern PDHG for solving large-scale LPs. Numerical experiments on both MIPLIB relaxations and Mittelmann’s benchmark set show that \cupdlpx consistently outperforms \texttt{cuPDLP}, achieving 2×–6× speedups across different tolerances and presolve settings. In particular, the C implementation provides further performance improvements over the Julia prototype. These results highlight the effectiveness of algorithmic and implementation enhancements in exploiting GPUs for high-performance linear programming.

\section*{Acknowledgement}
Haihao Lu is partially supported by AFOSR Grant No. FA9550-24-1-0051 and ONR Grant No. N000142412735. Zedong Peng is partially supported by ONR Grant No. N000142412735. Jinwen Yang is partially supported by AFOSR Grant No. FA9550-24-1-0051.

\bibliographystyle{amsplain}
\bibliography{ref-papers}

\end{document}